OpenStudy (anonymous):

Solve: √x+1+5=x My answer: x = 3 and x = 8

OpenStudy (anonymous):

\[\sqrt{x+1}+5=x\]Is this the problem

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

\[\sqrt{x+1}+5=x\]\[\sqrt{x+1}=x-5\]\[(\sqrt{x+1})^2=(x-5)^2\]\[x+1=x^2-10x+25\]\[0=x^2-11x+24\]\[0=(x-8)(x-3)\]\[x=3,8\]Perfect!!! :)

OpenStudy (anonymous):

oh my god that's exactly how i did my problem!!!!!!!!!!! thank you thank you thank you

OpenStudy (anonymous):

u should have more faith in yourself

OpenStudy (anonymous):

I'll try ;)